# Practical Secure Delegated Linear Algebra with Trapdoored Matrices

### v2

@@ Abstract
Most heavy computation occurs on servers owned by a second party. This reduces data privacy, resulting in interest in data-oblivious computation, which typically severely degrades performance. Secure and fast delegated computation is particularly important for linear algebra, which comprises a large fraction of total computation and is best run on highly specialized hardware often only accessible through the cloud.
We state the natural efficiency and security desiderata for fast and data-oblivious delegated linear algebra. We demonstrate the existence of trapdoored matrix families based on a LPN assumption, and provide a scheme for secure delegated matrix-matrix and matrix-vectors multiplication based on the existence of trapdoored matrices. We achieve sublinear overhead for the server, dramatically reduced computation cost for the client, and various practical advantages over previous algorithms.
Keywords:
Data Privacy, Data-Oblivious Computation, Delegation, Homomorphic Encryption, Cloud Computing, Algorithm Efficiency, Sublinear Overhead, LPN, Matrix Multiplication, Trapdoored Matrix
@@ 1 Introduction
Data-oblivious delegated computation – computation in which the server does not learn anything meaningful about the input data, thanks to cryptographic scrambling by the client – has been of theoretical and practical interest for some time [PF79, Ost90]. Arbitrary computation is quite difficult in this setting, requiring Fully Homomorphic Encryption (FHE) [RAD+78, Gen09] and currently incurring several orders of magnitude of overhead. As a result, much work on oblivious computation focuses on either advancing schemes for oblivious computation with more limited adversaries [MLS+13, ZE15], on developing and improving the runtime of FHE [FV12], or on Partially Homomorphic Encryption (PHE) [ElG85, GM84], which focuses on encryption that commutes with some operations.
A second paradigm, with different uses, has also emerged. Secure Multiparty Computation (MPC) [Yao86, GMW87], in which several non-colluding adversaries attempt computation on a union of datasets private to each server, has allowed for a variety of recent developments in hidden computation protocols. Here, a combination of informational and cryptographic techniques often offer efficiency improvements over corresponding PHE problems [MZ17, AFL+16, Kel20, KVH+21].
Much work in FHE and MPC has focused on lattice techniques. Learning With Errors (LWE), a survey of which is available in [Reg10], is the most well-known cryptographic problem in this area. Broadly speaking, the problem asks a user to distinguish a sequence of random values from a sequence of low-magnitude-noised inner products of known vectors $a_{i}$ against some unknown vector $s$. Learning Parity with Noise (LPN), a variant problem, instead asks that the noise be sparse.
In recent decades, an increasing fraction of computational power has been dedicated to matrix-vector and matrix-matrix multiplications. Specialized hardware to perform these operations is expensive and power-hungry, and is therefore unusually centralized. We often wish to allow low-power devices (such as personal computing hardware) to do large matrix-vector multiplications (in evaluating artificial neural networks, for instance) or to allow companies to rent computing power from centralized providers. However, input data must often be kept private (e.g. medical records, personal photos). We therefore desire schemes for delegated matrix-vector and matrix-matrix multiplication that are both secure and require minimal computational overhead – when even insecure computation is quite expensive, a constant-factor overhead is extremely costly. Techniques using exclusively operations native to modern GPUs/TPUs are preferred, as they benefit from extensive hardware and software optimization.
@@ 1.1 Our Contributions
We wish to conduct remote computations of the form $Av_{i}$ for some prespecified matrix $A$ and online vectors $v_{i}$, or matrix-matrix multiplications $AB$, while hiding $A,B,v_{i}$ from the remote server and minimizing local and total computational load. We state the natural efficiency and security desiderata, propose two trapdoored matrix objects, and provide an algorithm for the above linear-algebraic problems from such objects satisfying the aforementioned desiderata.
@@
We then provide an instantiation of such objects, and an associated algorithm, under a standard LPN [Kea98, HKL+12] hardness assumption. We also implement and benchmark our algorithm. It is lightweight and easy to implement, secure over any LPN-amenable ring or field, and uses only standard vector operations, rendering it highly amenable to compiler and hardware acceleration. As a result, it achieves both theoretically and practically small complexity blowup over computation in the clear while substantially reducing client computation.
1.2 Related Works
[Sot16] proposed a different version of trapdoored matrices, including a candidate construction. [^1] In this work, the definition of trapdoored matrices is not made exact, but weaker constructions (lacking pseudorandomness) are sufficient in their treated context, while stronger criteria are required in ours. We were unaware of this work in a previous version, and apologize for the omission.
@@ 1.2 Related Works
Several recent papers have focused directly on multiplication of matrices when one or both are encrypted, such as [JKLS18, BCH+24, LZ22, GQH+24]. [LZ22, BCH+24] in particular achieve efficient computation and decoding in practical regimes, but require expensive encoding. Many of these works also use LWE variants and incur a small accuracy error as a result.
@@
LPN-type problems were first discussed in [Kea98], and generalized to LPN in [HKL+12]. Best-known quantitative hardness result for LPN are discussed in [LWYY24].
@@
Oblivious linear function evaluation (OLE) was, to the best of our knowledge, first specialized as a problem of particular interest by [DGN+17], who provided improvements to the polynomial evaluation case of [NP99] (while we do not consider MPC or OLE directly, the following work represents some of the more advanced uses of LPN to date).[BCGI18, BCG+19, BCG+20b] demonstrated the usefulness of LPN for secure linear computation in MPC. In particular, [BCG+19] presented a MPC algorithm that can be turned into a vector-secure matrix-vector multiplication algorithm, but focused on goals corresponding to the MPC setting. [ADI+17, AK23] consider low-overhead MPC linear-algebra, including an algorithm obtaining constant-factor overhead for secure two-party arithmetic computation with honest adversaries. [CKR+20] used alternate techniques to achieve matrix multiplication with linear communication in an all-but-one-adversarial MPC framework.
@@
Secure matrix-vector multiplication is useful in various contexts, including evaluation of neural networks [MZ17, BEDK19, MWCB23].
@@
In a concurrent and independent work [VZ25], Vaikuntanathan and Zamir suggest a recursive construction of trapdoored matrices very similar to that of our Subsection 7 and use it to accelerate algorithms that use random matrices. Their construction can be used to slightly improve our server cost of matrix-matrix (removing the term in $\delta$) while approximately doubling client cost.
@@
A very recent follow-up work [BCH+25] has since provided alternate fast-matrix-vectors and Trapdoored-Matrix constructions using novel cryptographic assumptions, enabling significant storage and performance improvements in some domains. We expect further improved Trapdoored-Matrix constructions, with associated improvements in linear algebra delegation.
2 Fast Remote Computation
@@
  • The aggregate computation is very low (at most low-factor linear in, and ideally asymptotically equal to, the naive computational cost of the function).
  • The aggregate computation by the client is much less than that required to compute the function on its own (and ideally, nearly linear in input size).
  • The protocol is zero-knowledge for dishonest servers.
@@
For security, we use the standard definition:
@@
We say that a protocol is $(T(\cdot,\cdot),\epsilon(\cdot,\cdot),N(\cdot))$-secure if it has a security parameter $k$ such that, for all $n\geq N(k)$, the distributions of server views generated by any two client inputs are $(T(n,k),\epsilon(n,k))$-indistinguishable.
For efficiency, we compare to the unencrypted cost of the function:
Definition 2
We say that a protocol is $(c_{c}(n,k),c_{s}(n,k))$-remote-efficient if the client and the server expend $c_{c}(n,k)$ and $c_{s}(n,k)$ times the computation required to compute the function without encryption.
An ideal protocol would have $c_{c}(n,k)=0,c_{s}(n,k)=1$; an efficient protocol has $c_{c}(n,k)=o(1)$ and $c_{s}(n,k)=1+o(1)$
@@ 3 Three Problems and One Solution
Given $A\in R^{m\times n},B\in R^{n\times l}$ over a ring $R$, return $AB$.
@@
Note: a secure remote protocol for Matrix-Vectors yields a natural protocol for Matrix-Matrix (by multiplying one column of $B$ at a time), but that protocol does not enjoy the same remote efficiency as the Matrix-Vectors protocol (as the unencrypted complexity of Matrix-Matrix is lower than $n$-repeated Matrix-Vectors).
@@
Our results depend on a LPN hardness assumption, and have performance directly proportional to the strength of that assumption. To state performance concisely, let $\nu(\delta,\epsilon,\lambda)=\max\left(\frac{\lambda}{\lg|R|},\nu\right)$, where $R$ is the ring over which we compute and $\nu$ is the smallest value such that LPN in dimension $\frac{\nu}{\delta}$ with subspace dimension $\nu$ and noise rate $\left(\frac{\nu}{\delta}\right)^{\epsilon-1}$ has $\lambda$-bit security.
We then show the following:
@@
For any ring $R$ over which the LPN assumption holds, Algorithm 2 is a secure delegated Matrix-Matrices algorithm with parameters $\delta,\epsilon$ restricted to domains where the LPN assumption holds and security parameter $\lambda$. It guarantees:
@@
  • Preprocessing costs
    $O\left(n\left(n^{\epsilon}(m+n)+\nu(\delta,\epsilon,\lambda)(\delta n+m)\right)\right)$
@@
    $O\left(n^{\omega}\delta^{2\omega-3}+\min(m,\delta n)^{\omega-2}n\max(m,\delta n)\right)$
@@
    $O((m+n)l\nu(\delta,\epsilon,\lambda)+(m+n)n^{\epsilon}l)$
@@
  • Two-round preprocessing and one-round online steps.
  • Communication (both in preprocessing and in online steps) of $1+O\left(\delta\right)$ times input size.
In particular, in parameter regimes where $m=\Omega(n)$, the algorithm is
$\left(n^{\epsilon}+\frac{\nu(\delta,\epsilon,\lambda)}{n},1+O\left(\delta^{\omega-2}\right)\right)$
remote-efficient in both Matrix-Matrix and $\max(n^{\epsilon},\delta n^{\omega-2})$-steps-amortized Matrix-Vectors.
This specializes:
@@
Algorithm 2 restricts to a two-round [^3] A minor modification is required to parallelize independent server calls. secure delegated $n$-by-$n$ matrix-matrix protocol with parameter $\delta$ and client cost $O(n^{2}(n^{\epsilon}+\nu(\delta,\epsilon,\lambda)))$ and server cost beyond a single $n$-by-$n$ matrix multiplication of $O\left(\delta n^{\omega}\right)$.
@@
Algorithm 2 restricts to a two-preprocessing-round, one-round-per-online secure delegated $n$-by-$n$ matrix-vectors protocol with parameter $\delta$ and:
@@
  • Client preprocessing cost $O(n^{2}(n^{\epsilon}+\nu(\delta,\epsilon,\lambda)))$
@@ 4 Primitives for Fast Linear Algebra
Our approach to security is a delegation-oriented modification of Beaver’s triples [Bea92]: to secure the computation $AB_{i}$, generate pseudorandom matrices $A^{\prime}$ and $B^{\prime}_{i}$. Request the computation $(A+A^{\prime})(B_{i}+B^{\prime}_{i})$ from the server, and return $(A+A^{\prime})(B_{i}+B^{\prime}_{i})-A^{\prime}B_{i}-(A+A^{\prime})B^{\prime}_{i}$ (where the latter two terms are computed locally). While this is clearly secure by the same argument, it is not clearly efficient: it requires two local matrix multiplications. To improve efficiency, we must pick $A^{\prime}$ and $B^{\prime}_{i}$ to reduce the cost of these operations, inspiring the following cryptographic primitive generalizing an idea of [Sot16].
@@
A $(T(\cdot,\cdot),\epsilon(\cdot,\cdot),N(\cdot))$-secure, $C(\cdot,\cdot,\cdot)$-efficient Trapdoored-Matrix scheme is an algorithm that, given a ring $R$ and $m,n,l,k\geq N((m,n,R))\in\mathbb{Z}$, generates a $(T((m,n,R),k),\epsilon((m,n,R),k))$-pseudorandom $m\times n$ matrix $M$ over $R$ and $\tilde{O}(C(m,n))$ additional data such that, for any $N\in R^{n\times l}$, $MN$ can be computed in time $C(m,n,l)$. [^4] Note that this assumption is stronger than the fast-pseudorandom-matrix assumption of [ADI+17]: that assumption only requires that images under the matrix are random once noised. Many constructions satisfying that assumption (e.g. any sparse matrix) are insufficient here.555[Sot16] proposes a construction under a nonstandard security assumption. A construction of [BCG+20b] (sec. 10.3; the first construction with the Toeplitz/quasi-cyclic assumption) also fulfills these criteria, albeit with significantly stronger LPN assumptions and worse performance at equal security levels (as it was optimized for a substantially different task).
This scheme suffices for matrix-matrix multiplication. However, we wish also to hide matrix-vector multiplication, and therefore to develop pseudorandom trapdoored-vectors whose product with an arbitrary matrix $A$ can be computed in time much less than the naive product. This is impossible for one-shot instances, as the time to compute $Av$ for arbitrary $v$ is asymptotically equal to the time required to read $A$; therefore, we consider a further generalization of the above.
@@
A $(T(\cdot,\cdot),\epsilon(\cdot,\cdot),N(\cdot))$-secure, $C(\cdot,\cdot,\cdot)$-efficient Targeted-Trapdoored-Matrix scheme with $D(\cdot,\cdot,\cdot)$ preprocessing is an algorithm that, given a ring $R$, $m,n,l,k\geq N((m,n,R))\in\mathbb{Z}$, and a matrix $B\in R^{n\times l}$, preprocesses $B$ in time $D(\cdot,\cdot,\cdot)$, returning a $\tilde{O}(D(m,n))$-sized generator that generates $(T((m,n,R),k),\epsilon((m,n,R),k))$-pseudorandom $m\times n$ matrices $M_{i}$ and the associated value $M_{i}B$ in time $C(m,n,l)$.
Combined, these make the template scheme for Matrix-Matrix fast even when $B_{i}$ is a vector: use the Trapdoored-Matrix generator to generate $A^{\prime}$, and use the Targeted-Trapdoored-Matrix generator to generate the various $B^{\prime}_{i}$.
Assuming a large cache, we can obtain amortized Targeted-Trapdoored-Matrix from Trapdoored-Matrix: generate a trapdoored matrix $M$, compute $AM$, and then use blocks of $M$ as the $B^{\prime}_{i}$. Since we already need to store $A$, this space overhead is constant-factor.
We know of many families of matrices (Vandermonde matrices, Cauchy Matrices, discrete Chebyshev matrices) for which multiplication is $\tilde{O}(n)$. Since linear combinations of these or other fast-multiplication matrices could plausibly be pseudorandom, it seems natural to believe the following:
@@
There exist Trapdoored-Matrix constructions with preprocessing cost $\tilde{O}(n^{2})$ and matrix-vector multiplication cost $\tilde{O}(n)$.
The complexity of Targeted-Trapdoored-Matrix in the case where $m=1$ (i.e. Targeted-Trapdoored-Vectors) is less clear. The security assumption of Section 5 is commonly believed to hold for any $\delta,\epsilon>0$. Security of LPN in the dimension regime required for polylogarithmic overhead, however, is less plausible; if a polylogarithmic-overhead scheme exists, it will likely require another approach.
5 A LPN Hardness Assumption and Performance Scaling
Given a ring $R$, a dimension $n$, a subspace dimension $n_{1}$, and a noise rate $\mu$, the decisional LPN problem asks one to distinguish messages of the form $(L,Lr+s)$ from messages of the form $(L,u)$ where $L\overset{{\scriptscriptstyle\$}}{\leftarrow}R^{n\times n_{1}}$, $r\overset{{\scriptscriptstyle\$}}{\leftarrow}R^{n_{1}}$, $u\overset{{\scriptscriptstyle\$}}{\leftarrow}R^{n}$, and $s\leftarrow\mathcal{D}_{R,\mu}^{n}$ where $\mathcal{D}_{R,\mu}$ is the distribution that is 0 with probability $1-\mu$ and uniform over $R$ with probability $\mu$. A standard hardness assumption for LPN (as in [Ale03, BCGI18, YZ21]) is the following:

### v3

@@ Abstract
Most heavy computation occurs on servers owned by a second party. This reduces data privacy, resulting in interest in data-oblivious computation, which typically severely degrades performance. Secure and fast delegated computation is particularly important for linear algebra, which comprises a large fraction of total computation and is best run on highly specialized hardware often accessible only through the cloud.
We state the natural efficiency and security desiderata for fast and data-oblivious delegated linear algebra. We demonstrate the existence of Trapdoored-Matrix families based on an LPN assumption, and provide a scheme for secure delegated matrix-matrix and matrix-vector multiplication based on the existence of trapdoored matrices. We achieve sublinear overhead for the server, dramatically reduced computation for the client, and various practical advantages over previous protocols.
@@ 1 Introduction
Data-oblivious delegated computation – computation in which the server does not learn anything meaningful about the input data, thanks to cryptographic scrambling by the client – has been of theoretical and practical interest for some time [35, 34]. Arbitrary computation is quite difficult in this setting, requiring Fully Homomorphic Encryption (FHE) [37, 21] and currently incurring several orders of magnitude of overhead. As a result, much work on oblivious computation focuses on either advancing schemes for oblivious computation with more limited adversaries [31, 42], on developing and improving the runtime of FHE [18], or on Partially Homomorphic Encryption (PHE) [17, 23], which focuses on encryption that commutes with some operations.
A second paradigm, with different uses, has also emerged. Secure Multiparty Computation (MPC) [40, 22], in which several non-colluding adversaries compute on a union of datasets private to each server, has allowed for a variety of recent developments in hidden computation protocols. Here, a combination of informational and cryptographic techniques often offer efficiency improvements over corresponding PHE problems [33, 4, 27, 28].
Much work in FHE and MPC has focused on lattice techniques. Learning With Errors (LWE), a survey of which is available in [36], is the most well-known cryptographic problem in this area. Broadly speaking, the problem asks a user to distinguish a sequence of random values from a sequence of low-magnitude-noised inner products of known vectors $a_{i}$ against some unknown vector $s$. Learning Parity with Noise (LPN), a variant problem, instead asks that the noise be sparse.
In recent decades, an increasing fraction of computational power has been dedicated to matrix-vector and matrix-matrix multiplications. Specialized hardware to perform these operations is expensive and power-hungry, and is therefore unusually centralized. We often wish to allow low-power devices (such as personal computing hardware) to perform large matrix-vector multiplications (in evaluating artificial neural networks, for instance) or to allow companies to rent computing power from centralized providers. However, input data must often be kept private (e.g. medical records, personal photos). We therefore desire schemes for delegated matrix-vector and matrix-matrix multiplication that are both secure and highly efficient – when even insecure computation is quite expensive, a constant-factor overhead is extremely costly. Techniques using exclusively operations native to modern GPUs/TPUs are preferred, as they benefit from extensive hardware and software optimization.
@@ 1.1 Our Contributions
We wish to conduct delegated computations of the form $Av_{i}$ for some prespecified matrix $A$ and online vectors $v_{i}$, or matrix-matrix multiplications $AB$, while hiding $A,B,v_{i}$ from the server and minimizing client and total computational load. We propose a protocol for these problems using two novel cryptographic primitives: trapdoored matrices, matrices which are cryptographically pseudorandom but which can be multiplied by other matrices quickly using trapdoor information, and targeted trapdoored matrices, a slight generalization of trapdoored matrices that amortizes certain costs. At a high level, our protocol uses trapdoored matrices to additively mask operands before requesting their product from the server, and then uses the trapdoor property to quickly calculate the product of the unmasked operands from the product of the masked operands.
@@
We then provide constructions of trapdoored objects, and an associated protocol, under a standard LPN hardness assumption. We also implement and benchmark our protocol. It is lightweight and easy to implement, secure over any LPN-friendly ring or field, and uses only standard matrix and vector operations, rendering it highly amenable to compiler and hardware acceleration. As a result, it achieves both theoretically and practically small complexity blowup over computation in the clear while substantially reducing client computation.
1.2 Related Work
[38] [^1] We were unaware of this work in a previous version and apologize for the omission. proposed a different version of trapdoored matrices, including a candidate construction, while studying identity verification. In this work, the definition of trapdoored matrices is not made exact, but the properties required are far weaker. Instead of requiring trapdoored matrices to be pseudorandom, this work merely requires that it be challenging to compute information enabling quick multiplication from inspection of a randomly sampled trapdoor matrix.
@@ 1.2 Related Work
Several recent papers have focused directly on multiplication of matrices when one or both are encrypted, such as [25, 5, 30, 20]. [30, 5] in particular achieve efficient computation and decoding in practical parameter regimes, but require expensive (as compared to computation) encoding steps. Many of these works also use LWE variants and incur a small accuracy error as a result.
@@
LPN-type problems were first discussed in [8] and considered over finite fields in [24]. Algorithms for binary problems were considered in [26], and best-known quantitative hardness results for LPN over rings are discussed in [29].
@@
The link between oblivious linear function evaluation (OLE) and LPN was identified by [24] (while we do not consider MPC or OLE directly, this line of work represents some of the more advanced uses of LPN to date). [16, 9, 10, 12] further demonstrated the usefulness of LPN for secure linear computation in MPC. In particular, [10] presented a MPC protocol that can be turned into a vector-secure matrix-vector multiplication protocol, but focused on goals corresponding to the MPC setting. [2, 3] consider low-overhead MPC linear-algebra, including a protocol obtaining constant-factor overhead for secure two-party arithmetic computation with semi-honest adversaries. [13] used alternate techniques to achieve matrix multiplication with linear communication in an all-but-one-adversarial MPC framework.
@@
Secure matrix-vector multiplication is useful in various contexts, including evaluation of neural networks [33, 15, 32].
@@
In a concurrent and independent work [39], Vaikuntanathan and Zamir suggest a recursive construction of trapdoored matrices very similar to that of our Subsection 7 and use it to accelerate algorithms that use random matrices. Their construction can be used to slightly improve our server cost for matrix-matrix multiplication (removing the term in $\delta$) while approximately doubling client cost.
@@
A very recent follow-up work [7] has since provided alternate Trapdoored-Matrix and Targeted-Trapdoored-Matrix-like constructions using classical and novel cryptographic assumptions, enabling significant storage and performance improvements in some domains (see Section 8 for discussion). We expect further improved Trapdoored-Matrix constructions, with associated improvements in linear algebra delegation, to be developed in the future.
2 Fast Delegated Computation
@@
  • Total protocol computation is very low (at most low-factor linear in, and ideally asymptotically equal to, the naive computational cost of the function).
  • Total computation by the client is much less than that required to compute the function on its own (and ideally, nearly linear in input size).
  • The protocol is computational zero-knowledge for dishonest servers.
@@
For efficiency, we compare computational load to that of unencrypted delegation:
@@
We say that a client-server protocol in which the client receives data $A$ and is required to return prespecified $f(A)$ is $(c_{c},c_{s})$-efficient if the client and the server respectively expend $c_{c}$ and $c_{s}$ times their respective computational costs in the naive protocol in which the client sends $A$ to the server and the server computes and returns $f(A)$ to the client.
$c_{c}$ and $c_{s}$ will typically depend on parameters (e.g. problem size, security parameter). An ideal protocol would have $c_{c}=c_{s}=1$, an efficient protocol has $c_{c}$ small and $c_{s}=1+o(1)$, and a protocol beats local computation if $c_{c}$ is less than the cost of computing the result locally divided by the total input-output size.
@@ 3 Three Problems and One Solution
Given matrices $A\in R^{m\times n},B\in R^{n\times l}$ over a ring $R$, return $AB$.
@@
Note that a secure delegation protocol for Matrix-Vectors yields a natural protocol for Matrix-Matrix (by multiplying one column of $B$ at a time), but that protocol does not enjoy the same efficiency as the Matrix-Vectors protocol, as the unencrypted complexity of Matrix-Matrix is lower than $n$-repeated Matrix-Vectors.
@@
Our protocols for this problem have two phases: an initialization phase, in which the client and server interact for preprocessing purposes, and an online phase, in which the client and server interact after every online input.
Our results depend on an LPN hardness assumption (in Section 5), and have performance directly proportional to the strength of that assumption. To state performance concisely, let $\nu(\delta,\epsilon,\lambda)=\max\left(\frac{\lambda}{\lg|R|},\iota\right)$, where $R$ is the ring over which we compute and $\iota$ is the smallest value such that LPN in dimension $\frac{\iota}{\delta}$ with subspace dimension $\iota$ and noise rate $\left(\frac{\iota}{\delta}\right)^{\epsilon-1}$ obtains $(\lambda+\lg\lg n+1)$-bit-security CZK. We then show the following:
@@
For any ring $R$ over which the LPN assumption holds, Protocol 7 is a secure delegated Matrix-Matrices protocol. It guarantees, for any security parameter $\lambda$ and $\delta,\epsilon\in(0,1)$ such that the LPN assumption holds:
@@
  • Initialization costs
    $O\left(n(m+n)\left(n^{\epsilon}+\nu(\delta,\epsilon,\lambda)^{\omega-2}\right)\right)$
@@
    $O\left(n^{\omega}\delta^{\omega-1}+n\max(m,\delta n)\min(m,\delta n)^{\omega-2}\right)$
@@
    $O\left((m+n)l\left(n^{\epsilon}+\nu(\delta,\epsilon,\lambda)\right)\right)$
@@
  • Two-round initialization and one-round online steps.
  • Total communication of $1+O\left(\frac{\delta n}{m+l}\right)$ times input size.
In particular, in parameter regimes where $m=\Omega(n)$, the protocol is
$\left(n^{\epsilon}+\nu(\delta,\epsilon,\lambda),1+O\left(\delta^{\omega-2}\right)\right)$
efficient in both Matrix-Matrix and $\max(n^{\epsilon},(\delta n)^{\omega-2})$-steps-amortized Matrix-Vectors.
Two special cases are of particular interest.
@@
Protocol 7 restricts to a two-round [^3] A minor modification is required to parallelize independent server calls. secure delegated $n$-by-$n$ matrix-matrix protocol with client cost $O(n^{2}(n^{\epsilon}+\nu(\delta,\epsilon,\lambda)))$, and server cost beyond a single $n$-by-$n$ matrix multiplication of $O\left(\delta^{\omega-2}n^{\omega}\right)$.
@@
Protocol 7 restricts to a two-round-per-initialization, one-round-per-online-step secure delegated $n$-by-$n$ matrix-vectors protocol with:
@@
  • Client initialization cost $O(n^{2}(n^{\epsilon}+\nu(\delta,\epsilon,\lambda)^{\omega-2}))$
@@ 4 Primitives for Fast Linear Algebra
Our basic approach is a delegation-oriented modification of Beaver’s triples [6]: to secure the computation $AB_{i}$, generate pseudorandom matrices $A^{\prime}$ and $B^{\prime}_{i}$. Request the computation $(A+A^{\prime})(B_{i}+B^{\prime}_{i})$ from the server (sending the results of the sums and requesting their product), and return $(A+A^{\prime})(B_{i}+B^{\prime}_{i})-A^{\prime}B_{i}-(A+A^{\prime})B^{\prime}_{i}$ (where the latter two terms are computed locally). While this is clearly secure by the usual argument, it is not clearly efficient: it requires two local matrix multiplications. To improve efficiency, we must pick $A^{\prime}$ and $B^{\prime}_{i}$ to reduce the cost of these operations, inspiring the following cryptographic primitive generalizing an idea of [38].
@@
A Trapdoored-Matrix scheme is an algorithm that, given a ring $R$ and size $(m,n)$, generates a pseudorandom $m\times n$ matrix $M$ over $R$ and additional data such that, for any $N\in R^{n\times l}$, $MN$ can be computed more efficiently than via naive multiplication.444[38] proposes a weaker construction under a nonstandard security assumption. A construction of [12] (the first construction with the Toeplitz/quasi-cyclic assumption in Section 10.3) also fulfills these criteria, albeit with significantly stronger LPN assumptions and worse performance at equal security levels (as it was optimized for a substantially different task). A work concurrent to and independent of ours [39] and a work subsequent to ours [7] propose other constructions.
This scheme suffices for matrix-matrix multiplication. However, we also wish to delegate matrix-vector multiplication, and therefore to develop pseudorandom trapdoored vectors whose product with an arbitrary matrix $A$ can be computed in time much less than the naive product. This is impossible for one-shot instances, as the time to compute $Av$ for arbitrary $v$ is asymptotically equal to the time required to read $A$; therefore, we consider a further generalization of the above.
@@
A Targeted-Trapdoored-Matrix scheme is an algorithm that, given a ring $R$, size $(m,n)$, $l\geq N((m,n,R))\in\mathbb{Z}$, and a matrix $M\in R^{n\times l}$, preprocesses $M$ and returns an online generator that generates pseudorandom $m\times n$ matrices $B_{i}$ and the associated values $B_{i}M$ in less time than is required to draw a uniformly random $B_{i}$ and compute the products $B_{i}M$.
Combined, these make the template scheme for Matrix-Matrices fast even when $B_{i}$ is a vector: use the Trapdoored-Matrix generator to generate $A^{\prime}$, and use the Targeted-Trapdoored-Matrix generator to generate the various $B^{\prime}_{i}$.
At the cost of $O((n+m)tl)$ additional storage and of beginning the protocol with an $m\times n$-by-$n\times tl$ secured delegated matrix-matrix computation, we can obtain amortized Targeted-Trapdoored-Matrix for a $t$-matrix stream from Trapdoored-Matrix: generate a trapdoored matrix $M$, compute $AM$, and then use blocks of $M$ and $AM$ as the $B^{\prime}_{i}$ and $AB^{\prime}_{i}$, respectively.
We know of many families of matrices (e.g. Vandermonde, Cauchy, and discrete Chebyshev matrices) for which multiplication is $\tilde{O}(n)$. Since linear combinations of these or other matrices admitting fast multiplication could plausibly be pseudorandom, it seems natural to believe the following conjecture [^5] An update: a subsequent work, [7], provides conjectured-secure Trapdoored-Matrix constructions with $O(n)$ multiplication cost.:
@@
There exist Trapdoored-Matrix constructions with polynomial-time computational security, preprocessing cost $\tilde{O}(n^{2})$ and matrix-vector multiplication cost $\tilde{O}(n)$.
The complexity of Targeted-Trapdoored-Matrix in the case where $m=1$ (i.e. Targeted-Trapdoored-Vector) is less clear. The security assumption of Section 5 is commonly believed to hold for any $\delta,\epsilon>0$. Security of LPN in the dimension regime required for polylogarithmic overhead, however, is less plausible; if a polylogarithmic-overhead scheme exists, it will likely require another approach.
5 An LPN Hardness Assumption and Performance Scaling
Given a ring $R$, a dimension $n$, a subspace dimension $n_{1}$, and a noise rate $\mu$, the decisional LPN problem asks one to distinguish messages of the form $(L,LH+S)$ from messages of the form $(L,U)$ where $L\overset{{\scriptscriptstyle\mathdollar}}{\leftarrow}R^{n\times n_{1}}$, $H\overset{{\scriptscriptstyle\mathdollar}}{\leftarrow}R^{n_{1}\times\mathrm{\,poly}(n)}$, $U\overset{{\scriptscriptstyle\mathdollar}}{\leftarrow}R^{n\times\mathrm{\,poly}(n)}$, and $S\overset{{\scriptscriptstyle\mathdollar}}{\leftarrow}\mathcal{D}_{R,\mu}^{n\times\mathrm{\,poly}(n)}$ where $\mathcal{D}_{R,\mu}$ is the distribution that is 0 with probability $1-\mu$ and uniform over $R$ with probability $\mu$. A standard hardness assumption for LPN (as in [9, 41]) is the following: